OpenStudy (anonymous):

How do I find this? Find the inverse of g(x) is the inverse a function? g(x)=-3x^2+7 thank youuu.

OpenStudy (anonymous):

y = -3x^2 + 7 y - 7 = -3x^2 (y-7)/-3 = x^2 sqrt(y-7)/-3 = x x^2 fails the "horizontal line test" so technically it isn't invertible, but you could say the inverse is +/- sqrt(y-7)/-3